# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

### v1

@@ 4.2 NLHVT Case
We will consider a simple CGLMP Bell inequality involving
@@
where the $\Delta_{ij}$ are all integers and it is understood that $+\Delta_{ij}$ means $+\Delta_{ij}(\mathop{\mathrm{m}od}d)$.
@@
As in the LHVT case, there are two observables $\Omega_{A1},\Omega_{A2}$ for sub-system $A$ and two observables $\Omega_{B1},\Omega_{B2}$ for sub-system $B$. The eigenvectors for all the observables are as for the LHVT case. We will also consider quantum probabilities for the same maximally entangled state that was treated in the LHVT case.
@@
First we construct the NLHVT expression for $S$. We have
@@
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,\delta_{l,k+\Delta_{12}}\\ +\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\,\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\,\delta_{l,k+\Delta_{21}}\end{array}\right\}$ (19)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}\,\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,\delta_{l,k+\Delta_{12}}\,\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\,\delta_{k,m+\Delta_{22}}\,\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\,\delta_{m,j+\Delta_{21}}\,\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\end{array}\right\}$ (24)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(\delta_{j,l+\Delta_{11}}+\delta_{k,m+\Delta_{22}}\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(\delta_{l,k+\Delta_{12}}+\delta_{m,j+\Delta 1}\right)\end{array}\right\}$ (27)
The correspondiing expression in LHVT is
@@
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)\\ \times(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta 1})\end{array}\right\}$ (31)
5 DETERMINING NLHVT BELL INEQUALITIES
@@
First we consider the conditions involving the Kroneka deltas involved in the simple CGLMP inequality and the matrix approach used to treat them. .
@@ 5.1 Conditions and Basic Matrix Approach
To derive these inequalities we must consider the conditions where the various Kronecka deltas are equal to unity. These conditions are
@@
We will consider cases where; Case 1 - all of $C1,C2,C3,C4$ are true (T); Case 2 - three of $C1,C2,C3,C4$ are true and one is false (F); Case 3 - two of $C1,C2,C3,C4$ are true and two are false; Case 4 - one of $C1,C2,C3,C4$ is true and three are false. For these various cases the values of $\delta_{11,22}=\left(\delta_{j,l+\Delta_{11}}+\delta_{k,m+\Delta_{22}}\right)$ and $\delta_{12,21}=\left(\delta_{l,k+\Delta_{12}}+\delta_{m,j+\Delta 1}\right)$ are determined in constructing the Bell inequality.
It turns out that we must consider matrix equations of the form
@@
where $M$ is a square matrix and $\mathbf{v}$ is a column vector whose elements are some (or all) of the integers $j,k,l,m$ and $\mathbf{\Delta}$ is a column vector whose elements are sums of the various $\Delta_{ij}$. We will use the result that if the determinent of $M$ is zero ($|M|=0$) then there is only a possible solution for $\mathbf{v}$ if all the elements of the corresponding $\mathbf{\Delta}$ are zero. This will then lead to a set of requirements to lead to Bell inequalities: $S=4$, $S\leq 3$, $S\leq 2$, $S\leq 1$ for the Cases 1, 2, 3, 4.
@@ 5.2 Requirements for Bell Inequality Tests
These are derived in Appendix B. The resulting requirements are as follows.
@@ 5.2.1 Case1 - S4absent4$\leq$ 4≤ 4
The requirements are
@@ 5.2.2 Case 2 - S $\leq 3$
The requirements are
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (36)
If this requirement is satisfied the Bell inequality $S\leq 3$ would be guaranteed. Note that if this requirement is satisfied then that for $S\leq 4$ is not.  However those for $S\leq 2,S\leq 1$ could still be satisfied.
@@ 5.2.3 Case 3 - S $\leq 2$
Since we wish to guarantee that the Bell inequality $S\leq 2$ would occur, we can be certain of this by requiring that all the distinct inequalities in the last equation are satisfied. Accordingly, the $And/Or$ are replaced by $And$ and overall we have
$\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{12}\neq 0\quad And\quad\Delta_{21}\neq 0\quad And\quad\Delta_{22}\neq 0$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{11}\neq 0\quad\Delta_{12}+\Delta_{22}\neq 0\quad And\quad\Delta{}_{21}+\Delta_{22}\neq 0$
$\Delta_{11}+\Delta_{21}+\Delta_{22}$$\neq$ $0\quad And\quad\Delta_{22}+\Delta_{12}+\Delta_{11}\neq 0$
$\Delta_{12}+\Delta_{11}+\Delta_{21}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0$
Note that if these requirements are satisfied then that for $S\leq 3$ could still be satisfied. For example as $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}+\Delta_{22}\neq 0$, we still could have $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ which is the $S\leq 3$ requirement. However,  those for $S\leq 4$ are not consistent with those here for $S$ $\leq 2$..
@@ 5.2.4 Case 4 - S $\leq 1$
Since we wish to guarantee that the Bell inequality $S\leq 1$ would occur, we can be certain of this by requiring that all the distinct inequalities in the last equation are satisfied. Accordingly, the $And/Or$ are replaced by $And$ and overall we have
$\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{12}\neq 0\quad And\quad\Delta_{21}\neq 0\quad And\quad\Delta_{22}\neq 0$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{11}\neq 0\quad\Delta_{12}+\Delta_{22}\neq 0\quad And\quad\Delta_{21}+\Delta_{22}\neq 0$
Note that if these requirements are satisfied then those for $S\leq 2,S\leq 3$ could still be satisfied. For example as $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}\neq 0$, we still could have $\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ which is a $S\leq 2$ requirement. Similarly if $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}+\Delta_{22}\neq 0$ we could still have $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$, which is the $S\leq 3$ requirement. All the other $S\leq 2$ could be obtained similarly. However those for $S\leq 4$ are not consistent with those here for S $\leq 1$. .

### v2

@@ 3.4 NLHVT Case
We now consider a more general CGLMP Bell inequality involving
@@
where the outcome shifts $\Delta_{ij}$ are all integers and it is understood that $+\Delta_{ij}$ means $+\Delta_{ij}(\mathop{\mathrm{m}od}d)$.
@@
As in the LHVT case, there are two observables $\Omega_{A1},\Omega_{A2}$ for sub-system $A$ and two observables $\Omega_{B1},\Omega_{B2}$ for sub-system $B$. The eigenvectors for all the observables are (26), (LABEL:Eq.EigenvectorsObservB)  as for the LHVT case. We will also consider quantum probabilities for the same maximally entangled state (25) that was treated in the LHVT case. However,  we now use the NLHVT expressions for the probabilities and no longer invoke locality.
@@
First we construct the NLHVT expression for $S$. We have using the marginal probabilities in (LABEL:Eq.MargProb2)
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})\,\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1})\,\delta_{l,k+\Delta_{12}}$ (35)
$+\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2})\,\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2})\,\delta_{m,j+\Delta_{21}}$
@@
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{j,l+\Delta_{11}}\\ +\mathop{\displaystyle\sum}\limits_{k,l}\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{l,k+\Delta_{12}}\\ +\mathop{\displaystyle\sum}\limits_{k,m}\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{k,m+\Delta_{22}}\\ +\mathop{\displaystyle\sum}\limits_{j,m}\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\,\delta_{m,,j+\Delta_{21}}\end{array}\right\}$
@@
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\\ \times\{\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}}\}\end{array}\right\}$ (38)
Clearly the quantity $\delta=\{\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}}\}$ can have values $4,3,2,1$ depending on how many of the Kronecka deltas are equal to $1$ rather than  00. We see that the CGLMP Bell inequality in  the NLHVT case will just depend on the value for $\delta$
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\times\delta\right\}$ (39)
$\leq$ $\delta$
using $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$ and $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)=1$. In the next Section we determine sufficiency requirements on the shifts $\Delta_{ij}$ for the cases $\delta=4,3,2,1$.
By comparison with (38) the correspondiing expression if LHVT applies is
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)\\ \times(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}})\end{array}\right\}$ (42)
where here the locality based factorisation of $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ occurs. However the same sum of Kronecka deltas is present. In this LHVT version we see that we also have
$S\leq\delta$ (44)
using $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)=1$ as each factor is equal to $1$, for example $\sum_{j}P(\alpha_{j}|\Omega_{A1},\lambda)=1$.
This means that if the NLHVT is violated by quantum theory, then the LHVT will also be violated.
Note that the sum of the joint measurement probabilities where specific relationships occur between the measurement outcomes for the two sub-systems is in general less that the sum of the joint measurement probabilities for all such relationships. Thus $P(\Omega_{A1}{\small=}\overline{\Omega_{B1}+\Delta_{11}})=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}\leq\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)=P(\Omega_{A1},\Omega_{B1})=1$., where $P(\Omega_{A1},\Omega_{B1})$ is the overall probability for all outcomes of joint measurements for the choice $\Omega_{A1}$ and $\Omega_{B1}$. Similar considerations apply to $P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}\Delta_{12}})$, $P(\Omega_{A2}{\small=}\overline{\Omega_{B2}+\Delta_{22}})$ and $P(\Omega_{B2}{\small=}\overline{\Omega_{A1}+\Delta_{21})\text{.}}$.
4 REQUIREMENTS FOR NLHVT BELL INEQUALITIES
@@
First we consider the conditions involving the Kroneka deltas involved in the simple CGLMP inequality and the matrix approach used to treat them. Since the requirements only involve considering the situations where the Kronecka sum $\delta$ is given by $4,3,2$ or $1$, it follows that the requirements will be the same for both the LHVT and NLHVT cases.
@@ 4.1 Conditions and Basic Matrix Approach
To derive the sufficiency requirements on the outcome shifts $\Delta_{ij}$ for the NLHVT Bell Inequalities (39) we must consider the conditions where the various Kronecka deltas are equal to unity. These conditions are
@@
We will consider cases where; Case 1 - all of $C1,C2,C3,C4$ are true (T); Case 2 - three of $C1,C2,C3,C4$ are true and one is false (F); Case 3 - two of $C1,C2,C3,C4$ are true and two are false; Case 4 - one of $C1,C2,C3,C4$ is true and three are false. For these four Cases, there are a number of Sub-Cases depending on which of $C1,C2,C3,C4$ are true and which are false - one for Case 1, four for Case 2, six for Case 3 and four for Case 4. For the four Cases the value of $\delta=$ $(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}})$ are the same for every Sub-Case, namely $\delta=4,3,2,1$. for Cases $1,2,3,4$ respectively. As described in Sect.3.4 for these four Cases the values of $\delta$ are used to determine the Bell inequality $S\leq\delta$ (39) using Eq. (38) for $S$ (defined in Eq. (30)). However, the requirements on the shifts $\Delta_{ij}$ for each Sub-Case will be different, and these are set out below. Details are given in Appendix B. .
To derive these requirements it turns out that we must consider matrix equations of the form
@@
where $M$ is a square matrix and $\mathbf{v}$ is a column vector whose elements are some (or all) of the integers $j,k,l,m$ and $\mathbf{\Delta}$ is a column vector whose elements are sums of the various $\Delta_{ij}$. We will use the result that if the determinent of $M$ is zero ($|M|=0$) then there is only a possible solution for $\mathbf{v}$ if all the elements of the corresponding $\mathbf{\Delta}$ are zero. This will then lead to a set of requirements on the shifts $\Delta_{ij\text{ }}$to lead to Bell inequalities: $S\leq 4$, $S\leq 3$, $S\leq 2$, $S\leq 1$ for each Sub-Case of the Cases 1, 2, 3, 4. Note that we merely have to have one element of $\Delta$ to be non-zero to show there is no solution for $\mathbf{v}$.
The key point is that the CGLMP Bell Inequalities are of the form $S\leq\delta$, so they only depend on the number of the conditions $C1,C2,C3,C4$ that are satisfied. The basic strategy for each Sub Case is to take each condition that is required to be False, then substitute the conditions that are required to be True into these to arrive at the matrix equation (46) for the remaining indices from $j,k,l,m$ and then identify the column vector $\Delta$ for this  Sub-Case. We then show that the determinent $|M|$ is zero. As then there is to be no solution for the conditions that are False, we then obtain the requirement that at least one of  the elements of $\Delta$ is non-zero. Hence the requirement for each Sub-Case will be a list of And/Or statements. Each statement is a set of And/Or inequalities (or an equality in the Case of $S\leq 4$) arranged in rows for the various Sub-Cases, and involving the outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$. If at least one of these these inequalities are satisfied in every row, then we have obtained sufficiency requirements to show $S\leq\delta$ for the Case involved. Note that we do not obtain necessity requirements, as a particular set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ may provide sufficiency requirements for more than one of $S\leq 3$, $S\leq 2$ and  $S\leq 1$.  If more than one is satisfied then the overall NLHVT prediction is $S$leq the smallest of $3,2,1$ satisfied. This is logical since if for the outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ we can show that $S\leq 3$, $S\leq 2$ and $S\leq 1$ then overall we have $S\leq 1$.  The Case of $S\leq 4$ is an  exception, as we will see. Sufficiency requirements for $S\leq 4$ are incompatible with those for $S\leq 3$, $S\leq 2$ and  $S\leq 1$.
@@ 4.2 Requirements for Bell Inequality Tests
These Sub-Case requirements are derived in Appendix B for all Sub-Cases of the four Cases. These are summarised below in Eqs. (47), (48) , (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases). The resulting requirements for the four Cases is based on the distinct Sub-Case requirements and must cover all Sub-Cases to constitute sufficient (rather than necessary) requirements for concluding that according to our general NLHVT we have $S$ satisfying $S\leq 4$, $S\leq 3$, $S\leq 2$ or $S\leq 1$. These requirements are set out below in Eqs. (47), (48), (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases) respectively. In deriving these results we have used $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$ and $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)=1$.
@@ 4.2.1 Case1 - S4absent4$\leq$ 4≤ 4
The requirements for the single Sub-Case are
$\Delta_{11}=0\quad And\quad\Delta_{12}=0\quad And\quad\Delta_{22}=0\quad And\quad\Delta_{21}=0$ (47)
@@ 4.2.2 Case 2 - S $\leq 3$
The requirements for all four Sub-Cases are the same
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Cases\text{ }a,b,c,d$ (48)
If this requirement applies the Bell inequality $S\leq 3$ would be satisfied.
@@ 4.2.3 Case 3 - S $\leq 2$
Since we wish to guarantee that the Bell inequality $S\leq 2$ would occur irrespective of the Sub-Case, we can be certain of this by checking for a given set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ that at least one of the distinct inequalities in each row of the last equation is satisfied.  For exanple, the requirements $\Delta_{11}\neq 0,\Delta_{12}\neq 0,\Delta_{22}\neq 0$ and $\Delta_{21}\neq 0$ would satisfy the requirements for Sub-Cases $(a),(c),(d)$ and $(f)$ but not those for Sub-Cases $(b)$ and $(e)$, so with these four requirements alone we cannot conclude that $S\leq 2$.
@@ 4.2.4 Case 4 - S $\leq 1$
Since we wish to guarantee that the Bell inequality $S\leq 1$ would occur irrespective of the sub-case, we can be certain of this by checking for a given set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ that at least one of  the distinct inequalities in each row of the last equation are satisfied.
4.3 Example
As an example of what happens for a particular choice of outcome shifts we choose the case $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(1,-2,1,-2).$ It is obvious that the requirement for $S\leq 4$ is not satisfied.
We substitute the numerical values into Eqs. (48), (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases) .
For $S\leq 3$ (all Sub-Case) we see that $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}=(-2)\neq 0$ as required.
For $S\leq 2$we have
$\Delta_{22}$ $=$ $(2)\neq 0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}=(-3)\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $(-2)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}=(0)\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$ $=$ $(0)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}=(-3)\neq 0\qquad Sub-Case\text{ }f$
so as there is at least one  inequality satisfied for each Sub-Case we see that the requirements for $S\leq 2$are satisfied.
For $S\leq 1$we have
$\Delta_{12}$ $=$ $(-2)\neq 0\quad And/Or\quad\Delta_{22}=(1)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}=(-1)\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{22}=(1)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{12}=(-1)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{12}=(-2)\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }d$
so as there is at least one  inequality satisfied for each Sub-Case we see that the requirements for $S\leq 1$are satisfied.
Thus, the requirements for $S\leq 3$, $S\leq 2$ and $S\leq 1$.are all satisfied and we can conclude overall that NLHVT predicts $S\leq 1$ for this choice of outcome shifts.